On January 1, 2021, Morton Sales Co. issued zero-coupon bonds with a face value of $7.0 million for cash. The bonds mature in 8 years and were issued at a price of $3,781,890. (FV of $1, PV of $1, FVA of $1, PVA of $1, FVAD of $1 and PVAD of $1) (Use appropriate factor(s) from the tables provided.) Required: What amount of interest expense on these bonds would Morton Sales Co. report in its 2021 income statement?
Answer #1

Present value factor = 3781890 / 7000000 = 0.5403

Discount rate/ interest rate = 8%

Interest expense report in 2021 income statement = 3781890*8% = $302551